What volume of 0.558 M
1
Expert's answer
2020-09-01T07:37:34-0400

This question is only part of some task.


General solution:

1st Case:

What volume of 0.558 M A contains X moles of A.

Where A is some substance, X is the amount of a given substance.


Molarity of A = Moles of A / Volume of A

Then,

Volume of A = Moles of A / Molarity of A

Volume of A = Moles of A / 0.558 M


2nd Case:

"What volume of 0.558 M" - this is the beginning of a more complicated task.

Such as: "What volume of 0.558 M HNO3 solution is required to react completely with 45.55 mL of 0.0515 M Ba(OH)2?


Then,

Ba(OH)2 + 2HNO3 = Ba(NO3)2 + 2H2O

Moles of Ba(OH)2 = (0.04555 L)(0.0515 mol/L) = 0.002346 mol

According to the equation: Moles of Ba(OH)2 = Moles of HNO3 / 2

Moles of HNO3 = 2 * Moles of Ba(OH)2 = 2 * (0.002346 mol) = 0.004692 mol

Volume of HNO3 = Moles of HNO3 / Molarity of HNO3

Volume of HNO3 = (0.004692 mol) / (0.558 M) = 0.0084 L = 8.4 mL

V(HNO3) = 8.4 mL


However, to answer this question correctly requires a full question.